OpenStudy (anonymous):

Simplify the expression (4x^2/6y^4)^3 A. 2x^6/3y^12 B. 8x^6/27y^12 C. 27x^5/38y^7 D. 2x^5/3y^7

OpenStudy (anonymous):

We have : 4x^2+3/6y^4+4

OpenStudy (anonymous):

and : 4=2^2 and : 6=2*3

OpenStudy (anonymous):

ok, so the now it would look like this: \[4^{3}x ^{6} /6^{3}y ^{12}\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):